A Complete Bibliography of Publications in the SIAM Journal on Applied Algebra and Geometry

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

17 October 2023 Version 1.06

Title word cross-reference

2 [LW22]. 3 [DKPR22, MSS22]. C^{2} [CGY17]. C^{3} [HL23]. j [Kos22]. C^{∗} [MMW21]. F_{q} [MSS22]. R^{3} [HL23]. Z^{n} [YMV21]. μ [SG17]. n [GO19a]. p [KL21]. q [Jan23]. X [CQU17].

-Action [MMW21]. -Adic [KL21]. -Bases [SG17]. -Fold [MSS22]. -Invariants [Kos22]. -Matroid [Jan23]. -Parameter [LW22]. -Rank [CQU17]. -Zeros [MSS22].

Abstract [Fri23]. Action [MMW21]. Actions [MKSV20]. Adapted [HHS21]. Adic [KL21]. Alexander [GK17]. Algebra [DLW21, MKSV20]. Algebraic [AP23, BKM21, CMS17, DEW20, DM22, FdW22, KMZ21, KRV20, LS19, MS23, Sch18a, Sta21, TGND21]. Algorithm [Mic19, Sta21]. Algorithms [AKRS21, CFW21, LMR23]. Alternatives [AM19]. Ambiguity [Edi19]. Analogues [ACLN21]. Anomalies [SG17]. Application [CGY17]. Applications [BDFHH20, Der18, DGHP23, KK17, MS19]. Approach [BYF^{+}19, CMS21, CI22]. Approaches [AC22, Ave19]. Approximation [AP23, BDFHH20]. Arbitrary [ACLN21, BK18]. Arithmetic [Kop17]. Arrangements [GL22]. Aspects [Sta21]. Associated [SG17]. Atomically [YMV21]. Attacks [CLS17]. Auctions [JKS22]. Augmented [CKMW21]. Automorphisms [IL23]. Autonomy [SR18, Sha19]. Auxetic [BS17]. Averages [NW23]. Axiomatic [CMS21].

Bad [SG17]. Barrier [ABGJ18]. Barycenter [CS22]. Base [SG17]. Based [BYF^{+}19, Sta21, Tra21, Pol22]. Bases [KM19, SG17]. Bayesian [BSST23, Nic23]. Bent [GL22]. Betti [LW22]. Bigraded [LW22]. Bijectivity [MHR19]. Bilinear [KK17]. Biomembrane [CGY17]. Birationally [CDM18]. Blum [Sta21]. Bodies [FE18]. Body [GO19a]. Border [CGJ19, LM17]. Bottleneck [DEW20]. Bound [DV21, FE18, HM22]. Boundary [FdW22, LM20]. Bounds [MRZ22]. Branching [BCHP18]. Breaking [BCGO17]. Buildings [Hir19].

Calibrated [Kil17]. Can [BK20, Shi22]. Carathéodory [LR23]. Cartan [ÖSS18]. Case [BMN22]. CAT [HH21]. Cauchy [El 20]. Cells [BK18]. Central [BMN22, DH23]. Certifying [Sau19]. Characteristic [Mic19]. Characterization [AD23, NFB19]. Chief [Stu17]. Chiral [PST22]. Chordal [CP17]. Circuit [DIdW17, MS23, Wan22]. Circular [For23]. Classes [KE22]. Clustering [CMS21]. Code [CGJ^{+}17]. Coded [FHGHK17]. Codes [ACLN21, BNRS19, CFS19, GR22, Jef20, Jef22, LS18b, MSS22]. Cohomology [YL17]. Colluding [FHGHK17]. Colored [CMMS23]. Combinatorial [Rav22]. Comeagre [HL23]. Common [GR22]. Communications [BDFHH20]. Comon [Shi18]. Compact [HL23]. Compatibility [BN20]. Complements [GR22]. Complete [HL23, MMW21]. Completion [KKKR17]. Complex [HN23, Shi22]. Complexes [BHS^{+}19, MHP18]. Complexity [DIP20]. Computation [LS19, YMV21]. Computer [LV20]. Computing [CS22, HH21, Hir19, KMZ21, LM20, LW22, MS23]. Condition [Bür17]. Conditions [KK17]. Cone [FdW22, Sch21]. Cones [IL23].

Configurations [DH23]. Conformal [CS22]. Conjecture [Shi18, Tra21]. Connection [BYF^{+}19]. Connection-Based [BYF^{+}19]. Connections [MKSV20]. Connectivity [CMS17]. Consistency [MHP18]. Constants [BN20]. Constrained [MHP18]. Constraints [KMZ21]. Containment [FNT17]. Contraction [ÖSS18]. Control [BCGO17]. Controls [MHP18]. Convergence [BMN22]. Convex [BKM21, CGJ^{+}17, El 20, FE18, HH21, Hir19, Sch18a, Sch21]. Convexity [CFS19]. Convolutional [KMMT22]. Corrigendum [Sha19]. Cost [BCGO17]. Counterexample [Shi18]. Counting [CDM18, CHHM18]. Coupling [CHHM18]. Cryo [YL17]. Cryo-Electron [YL17]. Curvature [BKM21]. Curves [CCW19, KRV20, Kos22, Sch18a]. Cycle [Oba18]. Cycles [CCW19].

D [DKPR22]. Data [TGND21]. Databases [FHGHK17]. Datasets [NW23]. Decidability [CFS19]. Decomposing [DKPR22]. Decomposition [BHS^{+}19, CQU17, LMR23]. Decompositions [LM17]. Deconvolution [KK17]. Degree [BMN22, DEW20, Hir19, LNRW23, MRW20, MMW21, RW17, SR18, Sha19]. Denoising [Der18]. Density [Kra23]. Derivative [IL23]. Design [BS17]. Determinant [GO19a]. Determinantal [BFMT21, BvDD^{+}17, Vol19]. Determinants [Hir19]. Determined [Tur22]. Diagrams [MP22]. Differential [Cra19, DGHP23, KRV20]. Dimension [MM17]. Dimensional [Edi19, FNT17]. Dimensions [KOT23]. Diophantine [ADOO17]. Dirac [KW20]. Directed [CMS21]. Directions [DJ23]. Discrete [AA19, HH21, Hir19]. Distance [ALR19, MRW20, MP22, NFB19].

Distinguishing [GL18]. Distributed [DLW21]. Distributions [AA19, OPW20]. DLP [Mic19]. Do [BSV17]. Doubly [Dym18]. Dowling [Rav22]. Dream [BSV17]. Drone [BK20]. DSOS [AM19]. Duality [GK17]. Dynamical [Cra19, DGHP23].

Edit [MP22]. Editor [Stu17]. Editor-in-Chief [Stu17]. Eigenconfigurations [Koz18]. Eigenschemes [BGV21]. Eigenvalues [Bre17, KOT23, Kra23]. Eight [KOT23]. Elder [CKMW21]. Elder-Rule-Staircodes [CKMW21]. Electrical [For23]. Electron [YL17]. Elimination [DGHP23]. Elliptic [CCW19, Kos22]. Embeddings [HSW18]. Encodings [Kos22]. Enumerating [KE22]. Equation [RS21]. Equations [ADOO17]. Equilibria [CDM18, CHHM18]. Equivalence [LS18a]. Equivariant [CNN20]. Erratum [Jef22]. Errors [CLS17]. Estimation [AKRS21, DM21, MHP18]. Euclidean [HSW18, Hir19, MRW20, NFB19]. Evasion [GK17]. Evolution [LRR20]. Exact [Dym18]. Exceed [Shi22]. Exceptional [HSW18]. Existence [PST22]. Expected [Bre17]. experiment [OPPS22]. Explicit [NFB19]. Exponential [MHR19]. Exponential/Generalized [MHR19]. Expressiveness [GL22]. Extension [BdSN17]. Extensions [ACLN21, KMZ21]. Extraneous [SG17].

Facet [DJ23]. Factorization [CKM^{+}17]. Factors [SG17]. False [Tra21]. Families [BFMT21, MHR19]. Fan [MS19]. Fano [IZ17]. Fewnomials [BETC19]. Fibers [BBCY20]. Fields [Mic19]. Finite [CMS17, FNT17, Mic19, Tra21]. Finite-Dimensional [FNT17]. Finiteness [DH23]. Fixed [DH23, DJ23]. Flat [KOT23]. Flexible [CGY17]. Fold [MSS22].

Forced [Ber22]. Forms [El 20]. Formulas [BFMT21]. Foundations [LRR20, MKSV20]. Frames [CMS17, CI22]. Free [BSV17]. Friendly [CCW19]. Fully [Koz18]. Functions [AA19, BDFHH20, BCGO17, DEFT22, OPW20].

Galois [ACLN21, DS20, DKPR22].

Galois/Monodromy [DKPR22]. Gaussian [AA19, Bre17, CMMS23]. General [CM20, Der18, JPRS23]. Generalized [El 20, JKS22, MHR19]. Generated [IL23, YMV21]. Generator [Oba18]. Generic [Ber22, SR18, Sha19]. Genetic [LRR20]. Genomic [ALR19]. Geometric [AC22, DIP20, NW23]. Geometries [Rav22]. Geometry [BKM21, BDLS23, Edi19, KKKR17, KM19, KMMT22, Kop17, KL21, LS19, LM17]. Global [NFB19]. Gordan [LR23]. Gough [CM20, HSW18]. Graded [BBCY20]. Gram [HHS21]. Graph [CGG^{+}18, Pol22]. Graph-based [Pol22]. Graphical [CMMS23]. Graphs [AD23, Kra23]. Group [HSW18]. Groups [DKPR22, LS19, ÖSS18]. Guided [KE22].

Hausdorff [LM20]. Hear [BK20]. Hierarchical [CMS21]. Higher [BdSN17, HKU17, KM19, Rav22]. Higher-Order [HKU17]. Higher-Weight [Rav22]. Hilbert [HN23, LV20]. Hilbert-Samuel [HN23]. Homogeneous [Las19]. Homology [AC22, BK23, HOST19, LW22, MKPGC19, Oba18]. Hulls [Sch18a]. Hyperbolic [Sau19]. Hyperbolicity [IL23]. Hyperplane [GL22].

Ideals [CP17, CMMS23, KW20].

Identifiability

[CQU17, DGHP23, KK17, OPPS22].

Improvements [Sta21]. Including [Sta21]. Inclusion [BN20]. Inclusions [Cra19]. Independence [RW17]. Index [CDM18].

Inequalities [Ave19, El 20]. Inference [ALR19]. Infinitesimal [Ber22]. Infinity [SG17]. Information [DS20, FHGHK17]. Injectivity [KK17]. Innovation [LRR20]. Integer [OPW20]. Integral [KL21]. Integrals [BSST23]. Integration [MKSV20]. Interior [ABGJ18]. Interpolation [BdSN17]. Intersecting [Bür17]. Intersection [CDM18]. Invariant [AKRS21, CDM18, CI22, MKSV20]. Invariants [HL23, Kos22, LS19]. Inverse [KK17]. Irrational [CBG^{+}23]. Irrationality [CKM^{+}17]. Irreducibility [CMS17]. Isometries [YMV21]. Isoperimetric [DLSS19]. Isotopy [KE22]. Isotropy [CI22]. Iterated [DEFT22]. Iterated-Sums [DEFT22].

Jewel [BN20]. Joint [PSW21].

Kalai [Sta21]. Kernel [AP23]. Khovanskii [KM19]. Kippenhahn [PSW21]. Koszul [BFMT21]. Koszul-Type [BFMT21]. Kronecker [MM17]. Kuramoto [CDM18, CHHM18].

Laman [CGG^{+}18]. Laplacian [KOT23]. Large [KOT23]. Lattice [LS18a]. Lattices [BK18, MP22, Rav22]. Leaf [GO19b]. Learning [BYF^{+}19, DJ23]. Lemmas [MS19]. Lie [MKSV20]. Likelihood [AKRS21, DM21, LNRW23, MMW21, RW17]. Line [BRST23]. Linear [ADOO17, Ave19, Bür17, GR22, KMZ21, KMMT22, RS21]. Links [HN23]. Local [CFS19, KW20]. Locating [CHHM18]. Locus [CM20]. Log [ABGJ18, ALR19]. Log-Barrier [ABGJ18]. Log-Det [ALR19]. Lower [DV21, FE18]. LWE [Sta21].

Makes [CGJ^{+}17]. Manifold [BYF^{+}19]. Manifolds [MKSV20]. Many [BSS18]. Maps [BBCY20, CQU17, DS20, MHR19]. Marginals [MM17]. Mathematical [LRR20]. Matrices [NFB19]. Matrix [Ave19, BN20, CI22, CKM^{+}17, DM21, GG21, GO19a, LM17]. Matroid [Jan23, MS23, Tra21]. Matroid-Based [Tra21]. Maximum [AKRS21, DM21, LNRW23, MMW21, RW17]. Maxout [MRZ22]. Measure [LM20]. Measurements [BN20]. Message [Stu17]. Method [BSS18, LS19]. Methods [ABGJ18, TGND21]. Metric [AAF18, ACLN21, BNRS19, CKMW21]. Metrics [CR20]. Microscopy [YL17]. Minimal [DKPR22, Kil17, KE22, LW22]. Minimization [NFB19]. Minkowski [LR23, MRZ22]. Mixed [BFMT21]. Mixtures [KW20, RW17]. Model [CDM18, CHHM18, JPRS23, OPPS22]. Modelling [CGY17]. Models [CMMS23, DM21, DGHP23, MM17, RS21, RW17]. Modules [BdSN17]. Moment [DS20, KW20]. Monodromy [DKPR22]. Monomials [HM22]. Monotone [BDLS23]. Morphisms [Jef20, Jef22]. Morse [MHP18]. Motion [CFW21, ÖSS18]. MRD [GR22]. Muller [ACLN21]. Multi [BBCY20, OPPS22, RS21]. Multi-experiment [OPPS22]. Multi-Graded [BBCY20]. Multi-Trek [RS21]. Multilabeled [MS19]. Multilinear [BFMT21, DLW21]. Multiparameter [HOST19]. Multiplication [LM17]. Multiplicative [CGJ19]. Multiplicities [HN23]. Multiplicity [DIP20]. Multiview [BRST23, MRW20].

nc [HH21]. nc-Rank [HH21]. Networks [CMS21, CP17, Cra19, DP17, For23, GL22, GL18, KMMT22, MRZ22, Nic23]. Neural [CFS19, CGJ^{+}17, GL22, Jef20, Jef22]. Non [Nic23]. Non-toric [Nic23]. Noncommutative [Vol19]. Nonnegative [CKM^{+}17, DIdW17, Wan22]. Nonnegativity [Sau19]. Norm [CMS17]. Normal [DM21]. Norms [BK18, Nie17].

Nuclear [Nie17]. Number [Bre17, BETC19, CGG^{+}18, LS18b, MRZ22]. Numbers [LW22, Rav22]. Numerical [LS19, MKSV20, PSW21].

Obstruction [CFS19]. Obstructions [DIP20]. Occurrence [DIP20]. ODEs [OPPS22]. One [CR20, CHHM18, Edi19, IL23, KKKR17]. One-Dimensional [Edi19]. onto [BBCY20]. Operator [FNT17]. Operators [BKM21]. Optimal [AAF18, Ave19, BCGO17, JKS22, KK17, Kos22, Oba18]. Optimization [AM19, Ave19, BMN22, DM22, HH21, Hir19, OPW20, Sau19]. Optimize [Pol22]. Orbit [LS18a, YMV21]. Orbits [DS20]. Order [HKU17, Sch21]. Ordinary [GLSV19]. Orientable [HL23]. Orthogonal [BBCY20]. Other [LM17].

Packings [FJM18]. Pairing [CCW19]. Pairing-Friendly [CCW19]. Parallel [BYF^{+}19]. Parameter [CR20, LW22, OPPS22]. Parametric [OPW20]. Parametrized [CFW21]. Partial [GO19b]. Partitioned [NW23]. Path [BMN22]. Paths [BDLS23]. Paulsen [CI22]. Perfect [AD23]. Periodic [BS17, KE22, Kra23]. Permanent [Shi21]. Permutahedra [JKS22]. Persistence [BdSN17, CR20, MP22]. Persistent [AC22, BK23, HOST19, LW22, MKPGC19, Oba18]. Perspectives [DM22]. Perturbation [HKU17]. Phase [Edi19]. Phaseless [GG21]. Phylogenetic [DP17, For23, GO19b, GL18]. Picard [LS19, LS18b]. Pivot [BDLS23]. Planar [For23]. Plane [Sch21]. Planning [CFW21]. Platform [CM20]. Platforms [HSW18]. Point [ABGJ18]. Points [SG17]. Polyhedral [BSS18, BDLS23]. Polynomial [ABGJ18, Ave19, CP17, CQU17, Cra19, KMZ21, LNRW23, Mic19, MHR19, Pol22, Sau19]. Polynomials [DIdW17, LMR23, Las19, MS23, MSS22, Mic19, Vol19, Wan22].

Polytopes [BCHP18, DJ23, LS18a]. Posets [KMZ21]. Positive [BSS18, FE18, GK17]. Positivstellensatz [DIdW17]. Preordered [Fri23]. Presentations [LW22]. Principles [BS17]. Private [FHGHK17]. Problem [CLS17, CI22]. Problems [BCGO17, DKPR22, KK17, Kil17]. Product [CGJ19, MM17, SG17]. Profiles [DLSS19]. Programming [BS21]. Projection [BBCY20]. Projections [DGHP23]. Projective [Bür17, FJM18]. Projectivization [Jan23]. Propagation [JPRS23]. Pursuit [GK17].

Quadrics [MMW21]. Quantifying

[LRR20]. Quantum

[BN20, CBG^{+}23, DS20, PSW21]. Quasi [Mic19]. Quasi-Polynomial [Mic19]. Quasisymmetric [DEFT22]. Queueing [Pol22]. Quiver [CI22, DM21].

Radial [CI22]. Random [BETC19].

Ranges [PSW21]. Rank

[ACLN21, BNRS19, CGJ19, CQU17,

CHHM18, FE18, GG21, HH21, HM22, IL23, KKKR17, KM19, LM17, Shi21]. Rank-Metric [ACLN21, BNRS19]. Rank-One [CHHM18, IL23, KKKR17]. Rate [BMN22]. Rational [BBCY20, SG17]. Rationality [CM20]. Reaction [Cra19]. Real [Bre17, BETC19, FJM18, Koz18, NFB19, Sch18a]. Realization [FNT17]. Realizations [CGG^{+}18]. Reconstruction [AAF18, CNN20, DKPR22, GLSV19]. Reconstructions [PST22]. Recovery [Dym18]. Reduction [Pol22]. Reed [ACLN21]. Regions [MRZ22]. Related [LS19, OPW20]. Relaxation [Dym18]. Relaxations [IL23]. ReLU [GL22]. Representation [BNRS19, El 20, Sch18a, Sch21]. Representations [BvDD^{+}17, DM21, GO19b, Vol19]. Representative [Oba18]. Requires [CKM^{+}17]. Resolutions [BSV17]. Resonant [Kop17]. Respect [BK18]. Reticulate [LRR20]. Retrieval [Edi19, FHGHK17]. Rigidity [Ber22, MS23]. Ring [Sta21, Sta21]. Ring-Based [Sta21]. Ring-LWE [Sta21]. RLWE [CLS17]. RNA [BCHP18]. Robust [CMS21]. Room [BK20]. Rossby [Kop17]. Rotations [Ber22]. Rule [CKMW21]. Rules [BDLS23].

Samuel [HN23]. Scaling [AKRS21]. Scheme [CGY17]. Schemes [IZ17, LV20]. Schwarz [El 20]. SDSOS [AM19]. SE [HL23]. Search [CLS17]. Second [Sch21]. Second-Order [Sch21]. Segrè [HSW18]. Selection [Mic19]. Semialgebraic [BK23, LM20]. Semidefinite [AM19, Ave19, BMN22, BS21, FE18, Sch18a]. Semifields [Fri23]. Semirings [Fri23]. Separation [RS21]. Sequences [ALR19]. Series [DEFT22]. Servers [FHGHK17]. Set [HL23]. Sets [BK23, GO19b, Las19, LM20, Sch21]. Shadows [Sch18b]. Shape [BK20]. Sharp [MRZ22]. Signal [Der18]. Signatures [DEFT22, KRV20]. Signomial [DM22]. Silhouettes [GLSV19]. Simplicial [CNN20]. Singular [BHS^{+}19, Der18, HKU17, Tur22]. Singularities [CM20, GLSV19]. Size [Ave19]. Slope [JPRS23]. Smale [MHP18]. Small [CLS17, LS18b, Mic19]. Solutions [ADOO17, BSS18]. Solving [Sta21]. Sonc [FdW22]. Sonc-Cone [FdW22]. Space [DP17, GO19b]. Spaces [CKMW21, FJM18, HH21]. Sparse [ADOO17, BSS18, BS21, LNRW23, MSS22]. Sparseness [GR22]. Spatial [DH23]. Special [HSW18]. Species [ALR19]. Spectrahedra [HHS21]. Spectrahedral [FNT17, Sch18b]. Sperner [MS19]. Sphere [CGY17]. Splines [DV21]. Split [For23]. Square [NFB19]. Squares [AM19, AD23, BSV17, BS21, El 20].

Stabilization [CR20]. Stable [Vol19]. Staircodes [CKMW21]. Stars [DV21]. States [PSW21]. Statistical [MHP18]. Statistics [MKPGC19]. Stewart [CM20, HSW18]. Stochastic [Dym18]. Storage [DLW21]. Stratifying [HOST19]. Strong [SG17]. Stronger [DIP20]. Strongly [ABGJ18]. Structural [DGHP23, RS21]. Structure [BCHP18]. Study [LRR20]. Subconfigurations [DH23]. Subdivision [CGY17]. Subgroups [YMV21]. Subject [KMZ21]. Sublevel [Las19]. Subrank [Shi22]. Subspace [Bür17]. Subspaces [GR22]. Sufficient [MKPGC19]. Sum [AM19, AD23, El 20]. Sums [BSV17, BS21, DEFT22, DIdW17, MRZ22]. Surfaces [BBCY20, GLSV19, HL23, KE22, LS18b, SG17]. Symmetric [NFB19, Nie17, Shi22]. Symmetrically [CMMS23]. Symmetries [Dym18]. Symmetrized [NFB19]. Symmetry [Ber22, BCGO17, HHS21, KE22, LM17, LR23, Pol22]. Symmetry-Forced [Ber22]. Synchronization [ÖSS18]. Systems [BFMT21, BSS18, Cra19, For23, FNT17, LNRW23].

Template [MHP18]. Tensor

[Bre17, BNRS19, CGJ19, KKKR17, Nie17, SG17, Shi22, TGND21]. Tensors [BGV21, Koz18, LMR23, LM17, Tur22]. Ternary [BGV21]. Tetrahedral [DV21]. Their [GLSV19, IL23, Tur22, Vol19]. Theorem [PSW21]. Theorems [LR23]. Theoretic [CI22]. Theory [AKRS21, Der18, DS20, DIP20, OPPS22, Pol22]. Theta [AA19, BDFHH20]. Tight [CMS17]. Tightest [Oba18]. Tilings [KE22]. Time [DEFT22, Mic19]. Topological [GL22, LRR20]. Topologically [MHP18]. Topology [CFW21]. Tori [KOT23]. Toric [BSST23, CMMS23, Cra19, IZ17, MSS22, Nic23]. Total [DLSS19]. Tractable [AM19].

Transcendental [LS19]. Transformations [FN18]. Translations [Ber22]. Transport [AAF18, BYF^{+}19]. Transversality [GL22]. Tree [ALR19, GO19b]. Trek [RS21]. Triads [FN18, Kop17]. Trifocal [Kil17]. Triply [KE22]. Trivariate [MSS22]. Tropical [DEFT22, KM19, MKPGC19]. Tuples [Tur22]. Two [LV20, PST22]. Type [BFMT21].

Unfolding [BYF^{+}19]. Uniform [BvDD^{+}17]. Unit [CMS17]. Using [ALR19, CDM18].

Valuation [Tra21]. Valuations [KM19]. Value [BHS^{+}19]. Values [Der18, HKU17]. Vanishing [CMMS23]. Variation [DLSS19]. Varieties [AP23, BSST23, BRST23, CMS17, DEW20, IZ17]. Variety [Bür17, Kil17, MRW20]. Varying [Bür17]. Vergleichsstellensätze [Fri23]. Versions [MS19]. Vertex [DV21]. Vertices [MRZ22]. Via [AAF18, AA19, DM21, DGHP23, HH21, Hir19, MHP18, ÖSS18, Sau19]. View [PST22]. Vision [LV20]. Voicing [FN18]. Volume [Las19, Oba18]. Volume-Optimal [Oba18]. Voronoi [BK18].

Walks [CBG^{+}23]. Waring [HM22, Shi21]. Wasserman [Sta21]. Wave [Kop17]. Weight [Rav22]. Weyl [LR23]. Whitney [Rav22]. Wildfire [JPRS23]. Wind [JPRS23]. Worst [BMN22]. Worst-Case [BMN22].

x [Shi21].

Zeros [BETC19, MSS22].

References

Agostini:2019:DGD

[AA19] Daniele Agostini and Carlos Améndola. Discrete Gaussian distributions via theta functions. SIAM Journal on Applied Algebra and Geometry, 3(1):1–30, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Adamaszek:2018:MRO

[AAF18] Michał Adamaszek, Henry Adams, and Florian Frick. Metric reconstruction via optimal transport. SIAM Journal on Applied Algebra and Geometry, 2 (4):597–619, ???? 2018. CO-DEN SJAABQ. ISSN 2470-6566.

Allamigeon:2018:LBI

[ABGJ18] Xavier Allamigeon, Pascal Benchimol, Stéphane Gaubert, and Michael Joswig. Log-barrier interior point methods are not strongly polynomial. SIAM Journal on Applied Algebra and Geometry, 2(1):140–178, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Adams:2022:GAP

[AC22] Henry Adams and Baris Coskunuzer. Geometric approaches to persistent homology. SIAM Journal on Applied Algebra and Geometry, 6(4):685–710, ???? 2022. CODEN SJAABQ. ISSN 2470-

                                                   6566. URL https://epubs.
                                                   siam.org/doi/10.1137/21M1422914.

Augot:2021:RMC

[ACLN21] Daniel Augot, Alain Couvreur, Julien Lavauzelle, and Alessandro Neri. Rankmetric codes over arbitrary Galois extensions and rank analogues of Reed–Muller codes. SIAM Journal on Applied Algebra and Geometry, 5(2):165– 199, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1348583.

Ahmadi:2023:SSC

[AD23] Amir Ali Ahmadi and Cemil Dibek. A sum of squares characterization of perfect graphs. SIAM Journal on Applied Algebra and Geometry, 7(4):685– 715, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/22M1530410.

Aliev:2017:SSL

[ADOO17] Iskander Aliev, Jesús A. De Loera, Timm Oertel, and Christopher O’Neill. Sparse solutions of linear Diophantine equations. SIAM Journal on Applied Algebra and Geometry, 1(1):239–253, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Amendola:2021:ITS

[AKRS21] Carlos Améndola, Kathlén Kohn, Philipp Reichenbach, and Anna Seigal. Invariant theory and scaling algorithms for maximum likelihood estimation. SIAM Journal on Applied Algebra and Geometry, 5(2):304– 337, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/20M1328932.

Allman:2019:STI

[ALR19] Elizabeth S. Allman, Colby Long, and John A. Rhodes. Species tree inference from genomic sequences using the log-Det distance. SIAM Journal on Applied Algebra and Geometry, 3(1):107–127, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Ahmadi:2019:DSO

[AM19] Amir Ali Ahmadi and Anirudha Majumdar. DSOS and SDSOS optimization: More tractable alternatives to sum of squares and semidefinite optimization. SIAM Journal on Applied Algebra and Geometry, 3(2):193–230, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Altschuler:2023:KAA

[AP23] Jason M. Altschuler and Pablo A. Parrilo. Kernel approximation on algebraic varieties. SIAM Journal on Applied Algebra and Geometry, 7(1):1– 28, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1425050.

Averkov:2019:OSL

[Ave19] Gennadiy Averkov. Optimal size of linear matrix inequalities in semidefinite approaches to polynomial optimization. SIAM Journal on Applied Algebra and Geometry, 3(1):128–151, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Botbol:2020:FMG

[BBCY20] Nicolás Botbol, Laurent Busé, Marc Chardin, and Fatmanur Yildirim. Fibers of multi-graded rational maps and orthogonal projection onto rational surfaces. SIAM Journal on Applied Algebra and Geometry, 4 (2):322–353, ???? 2020. CO-DEN SJAABQ. ISSN 2470-6566.

Bloch:2017:OCP

[BCGO17] Anthony M. Bloch, Leonardo J.

Colombo, Rohit Gupta, and Tomoki Ohsawa. Optimal control problems with symmetry breaking cost functions. SIAM Journal on Applied Algebra and Geometry, 1(1):626–646, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Barrera-Cruz:2018:SRB

[BCHP18] Fidel Barrera-Cruz, Christine Heitsch, and Svetlana Poznanović. On the structure of RNA branching polytopes. SIAM Journal on Applied Algebra and Geometry, 2(3):444–461, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Barreal:2020:ATF

[BDFHH20] Amaro Barreal, Mohamed Taoufiq Damir, Ragnar Freij-Hollanti, and Camilla Hollanti. An approximation of theta functions with applications to communications. SIAM Journal on Applied Algebra and Geometry, 4 (4):471–501, ???? 2020. CO-DEN SJAABQ. ISSN 2470-6566.

Black:2023:PGP

[BDLS23] Alexander E. Black, Jesús A. De Loera, Niklas Lütjeharms, and Raman Sanyal. The polyhedral geometry of pivot rules and monotone paths. SIAM Journal on Applied Algebra and Geometry, 7(3):623– 650, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/22M1475910.

Bubenik:2017:HIE

[BdSN17] Peter Bubenik, Vin de Silva, and Vidit Nanda. Higher interpolation and extension for persistence modules. SIAM Journal on Applied Algebra and Geometry, 1(1):272–284, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Bernstein:2022:GSF

[Ber22] Daniel Irving Bernstein. Generic symmetry-forced infinitesimal rigidity: Translations and rotations. SIAM Journal on Applied Algebra and Geometry, 6(2): 190–215, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1346961.

Burgisser:2019:NRZ

[BETC19] Peter Bürgisser, Alperen A. Ergür, and Josué Tonelli-Cueto.

On the number of real zeros of random fewnomials. SIAM Journal on Applied Algebra and Geometry, 3(4):721–732, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Bender:2021:KTD

[BFMT21] Matías R. Bender, Jean-Charles Faugère, Angelos Mantzaflaris, and Elias Tsigaridas. Koszultype determinantal formulas for families of mixed multilinear systems. SIAM Journal on Applied Algebra and Geometry, 5(4): 589–619, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1332190.

Beorchia:2021:ETT

[BGV21] Valentina Beorchia, Francesco Galuppi, and Lorenzo Venturello. Eigenschemes of ternary tensors. SIAM Journal on Applied Algebra and Geometry, 5 (4):620–650, ???? 2021. CO-DEN SJAABQ. ISSN 2470-

6566. URL https://epubs.
siam.org/doi/10.1137/20M1355410.

Brake:2019:SVD

[BHS^{+}19] Danielle A. Brake, Jonathan D.

Hauenstein, Frank-Olaf Schreyer, Andrew J. Sommese, and Michael E. Stillman. Singular value decomposition of complexes. SIAM Journal on Applied Algebra and Geometry, 3 (3):507–522, ???? 2019. CO-DEN SJAABQ. ISSN 2470-

Blomer:2018:VCL

[BK18] Johannes Blömer and Kathlén Kohn. Voronoi cells of lattices with respect to arbitrary norms. SIAM Journal on Applied Algebra and Geometry, 2(2):314–338, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Boutin:2020:DCH

[BK20] Mireille Boutin and Gregor Kemper. A drone can hear the shape of a room. SIAM Journal on Applied Algebra and Geometry, 4(1):123–140, ???? 2020. CODEN SJAABQ. ISSN 2470-6566.

Basu:2023:PHS

[BK23] Saugata Basu and Negin Karisani. Persistent homology of semialgebraic sets. SIAM Journal on Applied Algebra and Geometry, 7(3):651–684, ???? 2023. CO-DEN SJAABQ. ISSN 2470-

                                        6566. URL https://epubs.
                                        siam.org/doi/10.1137/22M1494415.

Bettiol:2021:CAG

[BKM21] Renato G. Bettiol, Mario Kummer, and Ricardo A. E. Mendes. Convex algebraic geometry of curvature operators. SIAM Journal on Applied Algebra and Geometry, 5(2):200– 228, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/20M1350777.

Basu:2022:CPS

[BMN22] Saugata Basu and Ali Mohammad-Nezhad. On the central path of semidefinite optimization: Degree and worst-case convergence rate. SIAM Journal on Applied Algebra and Geometry, 6(2): 299–318, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1419933.

Bluhm:2020:CQM

[BN20] Andreas Bluhm and Ion Nechita. Compatibility of quantum measurements and inclusion constants for the matrix jewel. SIAM Journal on Applied Algebra and Geometry, 4(2):255–296, ???? 2020. CODEN SJAABQ. ISSN 2470-6566.

Byrne:2019:TRR

[BNRS19] Eimear Byrne, Alessandro Neri, Alberto Ravagnani, and John Sheekey. Tensor representation of rank-metric codes. SIAM Journal on Applied Algebra and Geometry, 3(4):614–643, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Breiding:2017:ENE

[Bre17] Paul Breiding. The expected number of eigenvalues of a real Gaussian tensor. SIAM Journal on Applied Algebra and Geometry, 1(1):254–271, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Breiding:2023:LMV

[BRST23] Paul Breiding, Felix Rydell, Elima Shehu, and Angélica Torres. Line multiview varieties. SIAM Journal on Applied Algebra and Geometry, 7(2):470– 504, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/22M1482263.

Borcea:2017:NPA

[BS17] Ciprian S. Borcea and Ileana Streinu. New principles for auxetic periodic design. SIAM Journal on Applied Algebra and Geometry, 1(1):442–458, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Blekherman:2021:SSS

[BS21] Grigoriy Blekherman and Kevin Shu. Sums of squares and sparse semidefinite programming. SIAM Journal on Applied Algebra and Geometry, 5(4): 651–674, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/20M1376170.

Bihan:2018:PMS

[BSS18] Frédéric Bihan, Francisco Santos, and Pierre-Jean Spaenlehauer. A polyhedral method for sparse systems with many positive solutions. SIAM Journal on Applied Algebra and Geometry, 2(4):620–645, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Borinsky:2023:BIT

[BSST23] Michael Borinsky, Anna-Laura Sattelberger, Bernd Sturmfels, and Simon Telen. Bayesian integrals on toric varieties. SIAM Journal on Applied Algebra and Geometry, 7(1):77– 103, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/22M1490569.

Blekherman:2017:DSS

[BSV17] Grigoriy Blekherman, Rainer Sinn, and Mauricio Velasco. Do sums of squares dream of free resolutions? SIAM Journal on Applied Algebra and Geometry, 1(1):175–199, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Burgisser:2017:CIP

[Bür17] Peter Bürgisser. Condition of intersecting a projective variety with a varying linear subspace. SIAM Journal on Applied Algebra and Geometry, 1(1):111–125, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Boralevi:2017:UDR

[BvDD^{+}17] Ada Boralevi, Jasper van Doornmalen, Jan Draisma, Michiel E. Hochstenbach, and Bor Plestenjak. Uniform determinantal representations. SIAM Journal on Applied Algebra and Geometry, 1(1):415–441, ???? 2017. CO-DEN SJAABQ. ISSN 2470-6566.

Budninskiy:2019:PTU

[BYF^{+}19] Max Budninskiy, Gloria Yin, Leman Feng, Yiying Tong, and Mathieu Desbrun. Parallel transport unfolding: a connection-based manifold learning approach. SIAM Journal on Applied Algebra and Geometry, 3(2):266–291, ???? 2019.

CODEN SJAABQ. ISSN 2470-

6566.

Coutinho:2023:IQW

[CBG^{+}23] Gabriel Coutinho, Pedro Ferreira Baptista, Chris Godsil, Thomás Jung Spier, and Reinhard Werner. Irrational quantum walks. SIAM Journal on Applied Algebra and Geometry, 7(3):567–584, ???? 2023. CO-DEN SJAABQ. ISSN 2470-

                                        6566. URL https://epubs.
                                        siam.org/doi/10.1137/22M1521262.

Chiesa:2019:CPF

[CCW19] Alessandro Chiesa, Lynn Chua, and Matthew Weidner. On cycles of pairing-friendly elliptic curves. SIAM Journal on Applied Algebra and Geometry, 3 (2):175–192, ???? 2019. CO-DEN SJAABQ. ISSN 2470-6566.

Chen:2018:CEK

[CDM18] Tianran Chen, Robert Davis, and Dhagash Mehta. Counting equilibria of the Kuramoto model using birationally invariant intersection index. SIAM Journal on Applied Algebra and Geometry, 2(4):489–507, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Chen:2019:NCD

[CFS19] Aaron Chen, Florian Frick, and Anne Shiu. Neural codes, decidability, and a new local obstruction to convexity. SIAM Journal on Applied Algebra and Geometry, 3(1):44–66, ???? 2019.

CODEN SJAABQ. ISSN 2470-6566.

Cohen:2021:TPM

[CFW21] Daniel C. Cohen, Michael Farber, and Shmuel Weinberger. Topology of parametrized motion planning algorithms. SIAM Journal on Applied Algebra and Geometry, 5(2):229–249, ???? 2021. CODEN SJAABQ. ISSN 2470-6566. URL https://

epubs.siam.org/doi/10.1137/
20M1358505.

Capco:2018:NRL

[CGG^{+}18] Jose Capco, Matteo Gallet, Georg Grasegger, Christoph Koutschan, Niels Lubbes, and Josef Schicho. The number of realizations of a Laman graph. SIAM Journal on Applied Algebra and Geometry, 2(1):94–125, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Curto:2017:WMN

[CGJ^{+}17] Carina Curto, Elizabeth Gross, Jack Jeffries, Katherine Morrison, Mohamed Omar, Zvi Rosen, Anne Shiu, and Nora Youngs. What makes a neural code convex? SIAM Journal on Applied Algebra and Geometry, 1(1):222–238, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Christandl:2019:BRM

[CGJ19] Matthias Christandl, Fulvio Gesmundo, and Asger Kjærulff Jensen. Border rank is not multiplicative under the tensor product. SIAM Journal on Applied Algebra and Geometry, 3 (2):231–255, ???? 2019. CO-DEN SJAABQ. ISSN 2470-6566.

Chen:2017:FSS

[CGY17] Jingmin Chen, Sara Grundel, and Thomas P. Y. Yu. A flexible C^{2} subdivision scheme on the sphere: With application to biomembrane modelling. SIAM Journal on Applied Algebra and Geometry, 1(1):459–483, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Coss:2018:LCE

[CHHM18] Owen Coss, Jonathan D. Hauenstein, Hoon Hong, and Daniel K. Molzahn. Locating and counting equilibria of the Kuramoto model with rank-one coupling. SIAM Journal on Applied Algebra and Geometry, 2(1):45–71, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Chindris:2022:QIT

[CI22] Calin Chindris and Jasim Ismaeel. A quiver invariant theoretic approach to radial isotropy and the paulsen problem for matrix frames. SIAM Journal on Applied Algebra and Geometry, 6(4):536– 562, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M141470X.

Chistikov:2017:NMF

[CKM^{+}17] Dmitry Chistikov, Stefan Kiefer, Ines Marušić, Mahsa Shirmohammadi, and James Worrell. Nonnegative matrix factorization requires irrationality. SIAM Journal on Applied Algebra and Geometry, 1(1):285–307, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Cai:2021:ERS

[CKMW21] Chen Cai, Woojin Kim, Facundo Mémoli, and Yusu Wang. Elder-rule-staircodes for augmented metric spaces. SIAM Journal on Applied Algebra and Geometry, 5(3):417– 454, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1353605.

Chen:2017:ASR

[CLS17] Hao Chen, Kristin Lauter, and Katherine E. Stange. Attacks on the search RLWE problem with small errors. SIAM Journal on Applied Algebra and Geometry, 1(1):665–682, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Coste:2020:RLS

[CM20] Michel Coste and Seydou Moussa. Rationality of the locus of singularities of the general Gough–Stewart platform. SIAM Journal on Applied Algebra and Geometry, 4(3):401–421, ???? 2020. CODEN SJAABQ. ISSN 2470-6566.

Coons:2023:SCG

[CMMS23] Jane I. Coons, Aida Maraj, Pratik Misra, and Miruna-Stefana Sorea. Symmetrically colored Gaussian graphical models with toric vanishing ideals. SIAM Journal on Applied Algebra and Geometry, 7(1):133– 158, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1466943.

Cahill:2017:CIA

[CMS17] Jameson Cahill, Dustin G. Mixon, and Nate Strawn. Connectivity and irreducibility of algebraic varieties of finite unit norm tight frames. SIAM Journal on Applied Algebra and Geometry, 1(1):38–72, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Carlsson:2021:RHC

[CMS21] Gunnar Carlsson, Facundo Mémoli, and Santiago Segarra. Robust hierarchical clustering for directed networks: An axiomatic approach. SIAM Journal on Applied Algebra and Geometry, 5(4):675–700, ???? 2021. CODEN SJAABQ. ISSN 2470-6566. URL https://

                                        epubs.siam.org/doi/10.1137/
                                        20M1359201.

Carbone:2020:ESR

[CNN20] Lisa Carbone, Vidit Nanda, and Yusra Naqvi. Equivariant simplicial reconstruction. SIAM Journal on Applied Algebra and Geometry, 4(4):532–552, ???? 2020. CODEN SJAABQ. ISSN 2470-6566.

Cifuentes:2017:CNP

[CP17] Diego Cifuentes and Pablo A.

Parrilo. Chordal networks of polynomial ideals. SIAM Journal on Applied Algebra and Geometry, 1(1):73–110, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Comon:2017:IRD

[CQU17] Pierre Comon, Yang Qi, and Konstantin Usevich. Identifiability of an X-rank decomposition of polynomial maps. SIAM Journal on Applied Algebra and Geometry, 1(1):388–414, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Chacholski:2020:MSO

[CR20] Wojciech Chachólski and Henri Riihimäki. Metrics and stabilization in one parameter persistence. SIAM Journal on Applied Algebra and Geometry, 4 (1):69–98, ???? 2020. CODEN SJAABQ. ISSN 2470-6566.

Craciun:2019:PDS

[Cra19] Gheorghe Craciun. Polynomial dynamical systems, reaction networks, and toric differential inclusions. SIAM Journal on Applied Algebra and Geometry, 3 (1):87–106, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Cantarella:2022:CCB

[CS22] Jason Cantarella and Henrik Schumacher. Computing the conformal barycenter. SIAM Journal on Applied Algebra and Geometry, 6(3):503– 530, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1449282.

Diehl:2022:TTS

[DEFT22] Joscha Diehl, Kurusch Ebrahimi-Fard, and Nikolas Tapia. Tropical time series, iterated-sums signatures, and quasisymmetric functions. SIAM Journal on Applied Algebra and Geometry, 6 (4):563–599, ???? 2022. CO-DEN SJAABQ. ISSN 2470-

6566. URL https://epubs.
siam.org/doi/10.1137/20M1380041.

Derksen:2018:GTS

[Der18] Harm Derksen. A general theory of singular values with applications to signal denoising. SIAM Journal on Applied Algebra and Geometry, 2(4):535–596, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

DiRocco:2020:BDA

[DEW20] Sandra Di Rocco, David Eklund, and Madeleine Weinstein. The bottleneck degree of algebraic varieties. SIAM Journal on Applied Algebra and Geometry, 4(1):227–253, ???? 2020. CODEN SJAABQ. ISSN 2470-6566.

Dong:2023:DED

[DGHP23] Ruiwen Dong, Christian Goodbrake, Heather A. Harrington, and Gleb Pogudin. Differential elimination for dynamical models via projections with applications to structural identifiability. SIAM Journal on Applied Algebra and Geometry, 7(1): 194–235, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/22M1469067.

Deng:2023:FSC

[DH23] Yiyang Deng and Marshall Hampton. Finiteness of spatial central configurations with fixed subconfigurations. SIAM Journal on Applied Algebra and Geometry, 7(3):610– 622, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/22M1490788.

Dressler:2017:PSN

[DIdW17] Mareike Dressler, Sadik Iliman, and Timo de Wolff. A Positivstellensatz for sums of nonnegative circuit polynomials. SIAM Journal on Applied Algebra and Geometry, 1(1):536–555, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Dorfler:2020:GCT

[DIP20] Julian Dörfler, Christian Ikenmeyer, and Greta Panova. On geometric complexity theory: Multiplicity obstructions are stronger than occurrence obstructions. SIAM Journal on Applied Algebra and Geometry, 4(2):354–376, ???? 2020. CO-DEN SJAABQ. ISSN 2470-6566.

Dostert:2023:LPF

[DJ23] Maria Dostert and Katharina Jochemko. Learning polytopes with fixed facet directions.

SIAM Journal on Applied Algebra and Geometry, 7(2):440– 469, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/22M1481695.

Duff:2022:GMG

[DKPR22] Timothy Duff, Viktor Korotynskiy, Tomas Pajdla, and Margaret H. Regan. Galois/Monodromy groups for decomposing minimal problems in 3D reconstruction. SIAM Journal on Applied Algebra and Geometry, 6(4):740– 772, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1422872.

DeFord:2019:TVI

[DLSS19] Daryl DeFord, Hugo Lavenant, Zachary Schutzman, and Justin Solomon. Total variation isoperimetric profiles. SIAM Journal on Applied Algebra and Geometry, 3(4):585–613, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Duursma:2021:MAD

[DLW21] Iwan Duursma, Xiao Li, and Hsin-Po Wang. Multilinear algebra for distributed storage. SIAM Journal on Applied Algebra and Geometry, 5(3):552– 587, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/20M1346742.

Derksen:2021:MLE

[DM21] Harm Derksen and Visu Makam.

Maximum likelihood estimation for matrix normal models via quiver representations. SIAM Journal on Applied Algebra and Geometry, 5(2):338– 365, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1369348.

Dressler:2022:APS

[DM22] Mareike Dressler and Riley Murray. Algebraic perspectives on signomial optimization. SIAM Journal on Applied Algebra and Geometry, 6(4):650– 684, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1462568.

Devadoss:2017:SPN

[DP17] Satyan L. Devadoss and Samantha Petti. A space of phylogenetic networks. SIAM Journal on Applied Algebra and Geometry, 1(1):683–705, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Dixon:2020:MMG

[DS20] Kael Dixon and Simon Salamon.

Moment maps and Galois orbits in quantum information theory. SIAM Journal on Applied Algebra and Geometry, 4(4):502–531, ???? 2020. CODEN SJAABQ. ISSN 2470-6566.

DiPasquale:2021:LBS

[DV21] Michael DiPasquale and Nelly Villamizar. A lower bound for splines on tetrahedral vertex stars. SIAM Journal on Applied Algebra and Geometry, 5(2): 250–277, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/20M1341118.

Dym:2018:ERS

[Dym18] Nadav Dym. Exact recovery with symmetries for the doubly stochastic relaxation. SIAM Journal on Applied Algebra and Geometry, 2(3):462–488, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Edidin:2019:GAO

[Edi19] Dan Edidin. The geometry of ambiguity in one-dimensional phase retrieval. SIAM Journal on Applied Algebra and Geometry, 3(4):644–660, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

ElKhadir:2020:SSR

[El 20] Bachir El Khadir. On sum of squares representation of convex forms and generalized Cauchy– Schwarz inequalities. SIAM Journal on Applied Algebra and Geometry, 4(2):377–400, ???? 2020. CODEN SJAABQ. ISSN 2470-6566.

ForsgAard:2022:ABS

[FdW22] Jens Forsgård and Timo de Wolff. The algebraic boundary of the sonc-cone. SIAM Journal on Applied Algebra and Geometry, 6(3):468–502, ???? 2022. CO-DEN SJAABQ. ISSN 2470-

6566. URL https://epubs.
siam.org/doi/10.1137/20M1325484.

Fawzi:2018:LBP

[FE18] Hamza Fawzi and Mohab Safey El Din. A lower bound on the positive semidefinite rank of convex bodies. SIAM Journal on Applied Algebra and Geometry, 2(1):126–139, ???? 2018. CO-DEN SJAABQ. ISSN 2470-6566.

Freij-Hollanti:2017:PIR

[FHGHK17] Ragnar Freij-Hollanti, Oliver W.

Gnilke, Camilla Hollanti, and David A. Karpuk. Private information retrieval from coded databases with colluding servers. SIAM Journal on Applied Algebra and Geometry, 1(1):647–664, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Fickus:2018:PRP

[FJM18] Matthew Fickus, John Jasper, and Dustin G. Mixon. Packings in real projective spaces. SIAM Journal on Applied Algebra and Geometry, 2(3):377–409, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Fiore:2018:VTT

[FN18] Thomas M. Fiore and Thomas Noll. Voicing transformations of triads. SIAM Journal on Applied Algebra and Geometry, 2 (2):281–313, ???? 2018. CO-DEN SJAABQ. ISSN 2470-

Fritz:2017:SCO

[FNT17] Tobias Fritz, Tim Netzer, and Andreas Thom. Spectrahedral containment and operator systems with finite-dimensional realization. SIAM Journal on Applied Algebra and Geometry, 1 (1):556–574, ???? 2017. CO-DEN SJAABQ. ISSN 2470-6566.

Forcey:2023:CPE

[For23] Stefan Forcey. Circular planar electrical networks, split systems, and phylogenetic networks. SIAM Journal on Applied Algebra and Geometry, 7 (1):49–76, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/22M1473844.

Fritz:2023:AVP

[Fri23] Tobias Fritz. Abstract Vergleichsstellensätze for preordered semifields and semirings I. SIAM Journal on Applied Algebra and Geometry, 7(2):505– 547, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/22M1498413.

Goucha:2021:PRM

[GG21] António Pedro Goucha and João Gouveia. The phaseless rank of a matrix. SIAM Journal on Applied Algebra and Geometry, 5 (3):526–551, ???? 2021. CO-DEN SJAABQ. ISSN 2470-

                                        6566. URL https://epubs.
                                        siam.org/doi/10.1137/19M1289820.

Ghrist:2017:PAD

[GK17] Robert Ghrist and Sanjeevi Krishnan. Positive Alexander duality for pursuit and evasion. SIAM Journal on Applied Algebra and Geometry, 1(1):308–327, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Gross:2018:DPN

[GL18] Elizabeth Gross and Colby Long. Distinguishing phylogenetic networks. SIAM Journal on Applied Algebra and Geometry, 2(1):72–93, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Grigsby:2022:TBH

[GL22] J. Elisenda Grigsby and Kathryn Lindsey. On transversality of bent hyperplane arrangements and the topological expressiveness of ReLU neural networks. SIAM Journal on Applied Algebra and Geometry, 6(2):216– 242, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1368902.

Gallet:2019:RSO

[GLSV19] Matteo Gallet, Niels Lubbes, Josef Schicho, and Jan Vrsek. Reconstruction of surfaces with ordinary singularities from their silhouettes. SIAM Journal on Applied Algebra and Geometry, 3(3):472–506, ???? 2019. CO-DEN SJAABQ. ISSN 2470-

Grinberg:2019:BMD

[GO19a] Darij Grinberg and Peter J. Olver. The n body matrix and its determinant. SIAM Journal on Applied Algebra and Geometry, 3(1):67–86, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Grindstaff:2019:RPL

[GO19b] Gillian Grindstaff and Megan Owen. Representations of partial leaf sets in phylogenetic tree space. SIAM Journal on Applied Algebra and Geometry, 3 (4):691–720, ???? 2019. CO-DEN SJAABQ. ISSN 2470-6566.

Gruica:2022:CCL

[GR22] Anina Gruica and Alberto Ravagnani. Common complements of linear subspaces and the sparseness of MRD codes. SIAM Journal on Applied Algebra and Geometry, 6(2):79– 110, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1428947.

Hamada:2021:CNR

[HH21] Masaki Hamada and Hiroshi Hirai. Computing the ncrank via discrete convex optimization on CAT(0) spaces. SIAM Journal on Applied Algebra and Geometry, 5(3):455– 478, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M138836X.

Heaton:2021:SAG

[HHS21] Alexander Heaton, Serkan Hosten, and Isabelle Shankar. Symmetry adapted Gram spectrahedra. SIAM Journal on Applied Algebra and Geometry, 5 (1):140–164, ???? 2021. CO-DEN SJAABQ. ISSN 2470-

6566. URL https://epubs.
siam.org/doi/10.1137/20M133796X.

Hirai:2019:CDD

[Hir19] Hiroshi Hirai. Computing the degree of determinants via discrete convex optimization on Euclidean buildings. SIAM Journal on Applied Algebra and Geometry, 3(3):523–557, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Hackbusch:2017:PHO

[HKU17] Wolfgang Hackbusch, Daniel Kressner, and André Uschmajew. Perturbation of higherorder singular values. SIAM Journal on Applied Algebra and Geometry, 1(1):374–387, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Hayut:2023:CSI

[HL23] Yair Hayut and David Lehavi.

Complete SE(3) invariants for a comeagre set of C^{3} compact orientable surfaces in R^{3}. SIAM Journal on Applied Algebra and Geometry, 7(2):311– 344, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1445776.

Han:2022:NBW

[HM22] Kangjin Han and Hyunsuk Moon. A new bound for the Waring rank of monomials. SIAM Journal on Applied Algebra and Geometry, 6(3): 407–431, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1390736.

Helmer:2023:CLH

[HN23] Martin Helmer and Vidit Nanda. Complex links and Hilbert-samuel multiplicities. SIAM Journal on Applied Algebra and Geometry, 7(1):29– 48, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/22M1475533.

Harrington:2019:SMP

[HOST19] Heather A. Harrington, Nina Otter, Hal Schenck, and Ulrike Tillmann. Stratifying multiparameter persistent homology. SIAM Journal on Applied Algebra and Geometry, 3(3):439–471, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Hauenstein:2018:ESG

[HSW18] Jonathan D. Hauenstein, Samantha N. Sherman, and Charles W. Wampler. Exceptional Stewart– Gough platforms, Segrè embeddings, and the special Euclidean group. SIAM Journal on Applied Algebra and Geometry, 2 (1):179–205, ???? 2018. CO-DEN SJAABQ. ISSN 2470-6566.

Ito:2023:ARO

[IL23] Masaru Ito and Bruno F.

Lourenço. Automorphisms of rank-one generated hyperbolicity cones and their derivative relaxations. SIAM Journal on Applied Algebra and Geometry, 7 (1):236–263, ???? 2023. CO-DEN SJAABQ. ISSN 2470-

6566. URL https://epubs.
siam.org/doi/10.1137/22M1513964.

Ilten:2017:FST

[IZ17] Nathan Ilten and Alexandre Zotine. On Fano schemes of toric varieties. SIAM Journal on Applied Algebra and Geometry, 1 (1):152–174, ???? 2017. CO-DEN SJAABQ. ISSN 2470-6566.

Jany:2023:PMQ

[Jan23]  Benjamin Jany.  The pro-

jectivization matroid of a qmatroid. SIAM Journal on Applied Algebra and Geometry, 7 (2):386–413, ???? 2023. CO-DEN SJAABQ. ISSN 2470-

6566. URL https://epubs.
siam.org/doi/10.1137/22M1494567.

Jeffs:2020:MNC

[Jef20] R. Amzi Jeffs. Morphisms of neural codes. SIAM Journal on Applied Algebra and Geometry, 4(1):99–122, ???? 2020. CODEN SJAABQ. ISSN 2470-6566. See erratum [?, ?].

Jeffs:2022:EMN

[Jef22] R. Amzi Jeffs. Erratum for “Morphisms of Neural Codes”.

SIAM Journal on Applied Algebra and Geometry, 6(1):77– 78, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1450690.

See [Jef20].

Joswig:2022:GPO

[JKS22] Michael Joswig, Max Klimm, and Sylvain Spitz. Generalized permutahedra and optimal auctions. SIAM Journal on Applied Algebra and Geometry, 6(4): 711–739, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1441286.

Javaloyes:2023:GMW

[JPRS23] Miguel Ángel Javaloyes, Enrique Pendás-Recondo, and Miguel Sánchez. A general model for wildfire propagation with wind and slope. SIAM Journal on Applied Algebra and Geometry, 7(2):414– 439, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/22M1477866.

Kolbe:2022:EIC

[KE22] Benedikt Kolbe and Myfanwy E. Evans. Enumerating isotopy classes of tilings guided by the symmetry of triply periodic minimal surfaces. SIAM Journal on Applied Algebra and Geometry, 6(1):1–40, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1358943.

Kileel:2017:MPC

[Kil17] Joe Kileel. Minimal problems for the calibrated trifocal variety. SIAM Journal on Applied Algebra and Geometry, 1(1):575–598, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Kech:2017:OIC

[KK17] Michael Kech and Felix Krahmer. Optimal injectivity conditions for bilinear inverse problems with applications to identifiability of deconvolution problems. SIAM Journal on Applied Algebra and Geometry, 1 (1):20–37, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Kahle:2017:GRO

[KKKR17] Thomas Kahle, Kaie Kubjas, Mario Kummer, and Zvi Rosen. The geometry of rank-one tensor completion. SIAM Journal on Applied Algebra and Geometry, 1(1):200–221, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Kulkarni:2021:AIG

[KL21] Avinash Kulkarni and Antonio Lerario. p-adic integral geometry. SIAM Journal on Applied Algebra and Geometry, 5 (1):28–59, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/19M1284737.

Kaveh:2019:KBH

[KM19] Kiumars Kaveh and Christopher Manon. Khovanskii bases, higher rank valuations, and tropical geometry. SIAM Journal on Applied Algebra and Geometry, 3(2):292–336, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Kohn:2022:GLC

[KMMT22] Kathlén Kohn, Thomas Merkh, Guido Montúfar, and Matthew Trager. Geometry of linear convolutional networks. SIAM Journal on Applied Algebra and Geometry, 6(3):368– 406, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1441183.

Kepley:2021:CLE

[KMZ21] Shane Kepley, Konstantin Mischaikow, and Lun Zhang. Computing linear extensions for polynomial posets subject to algebraic constraints. SIAM Journal on Applied Algebra and Geometry, 5(2):388– 416, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1343208.

Kopp:2017:AGR

[Kop17] Gene S. Kopp. The arithmetic geometry of resonant Rossby wave triads. SIAM Journal on Applied Algebra and Geometry, 1(1):352–373, ???? 2017. CO-DEN SJAABQ. ISSN 2470-6566.

Koshelev:2022:OEE

[Kos22] Dmitrii Koshelev. Optimal encodings to elliptic curves of j-invariants 0, 1728. SIAM Journal on Applied Algebra and Geometry, 6(4):600–617, ???? 2022. CODEN SJAABQ. ISSN 2470-6566. URL https://

epubs.siam.org/doi/10.1137/
21M1441602.

Kao:2023:FTL

[KOT23] Chiu-Yen Kao, Braxton Osting, and Jackson C. Turner. Flat tori with large Laplacian eigenvalues in dimensions up to eight. SIAM Journal on Applied Algebra and Geometry, 7(1):172– 193, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/22M1478823.

Kozhasov:2018:FRE

[Koz18] Khazhgali Kozhasov. On fully real eigenconfigurations of tensors. SIAM Journal on Applied Algebra and Geometry, 2 (2):339–347, ???? 2018. CO-DEN SJAABQ. ISSN 2470-6566.

Kravaris:2023:DEP

[Kra23] Cosmas Kravaris. On the density of eigenvalues on periodic graphs. SIAM Journal on Applied Algebra and Geometry, 7 (3):585–609, ???? 2023. CO-DEN SJAABQ. ISSN 2470-

6566. URL https://epubs.
siam.org/doi/10.1137/22M1505268.

Kogan:2020:DSA

[KRV20] Irina A. Kogan, Michael Ruddy, and Cynthia Vinzant. Differential signatures of algebraic curves. SIAM Journal on Applied Algebra and Geometry, 4 (1):185–226, ???? 2020. CO-DEN SJAABQ. ISSN 2470-6566.

Koutsoumpelias:2020:MIL

[KW20] Alexandros Grosdos Koutsoumpelias and Markus Wageringel. Moment ideals of local Dirac mixtures. SIAM Journal on Applied Algebra and Geometry, 4(1):1–27, ???? 2020. CODEN SJAABQ. ISSN 2470-6566.

Lasserre:2019:VSS

[Las19] Jean B. Lasserre. Volume of sublevel sets of homogeneous polynomials. SIAM Journal on Applied Algebra and Geometry, 3 (2):372–389, ???? 2019. CO-DEN SJAABQ. ISSN 2470-6566.

Landsberg:2017:GBR

[LM17] J. M. Landsberg and Mateusz Michałek. On the geometry of border rank decompositions for matrix multiplication and other tensors with symmetry. SIAM Journal on Applied Algebra and Geometry, 1(1):2–19, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Lasserre:2020:CHB

[LM20] Jean B. Lasserre and Victor Magron. Computing the Hausdorff boundary measure of semialgebraic sets. SIAM Journal on Applied Algebra and Geometry, 4 (3):441–469, ???? 2020. CO-DEN SJAABQ. ISSN 2470-6566.

Laface:2023:DAT

[LMR23] Antonio Laface, Alex Massarenti, and Rick Rischter. Decomposition algorithms for tensors and polynomials. SIAM Journal on Applied Algebra and Geometry, 7(1):264–290, ???? 2023. CODEN SJAABQ. ISSN 2470-

6566. URL https://epubs.
siam.org/doi/10.1137/21M1453712.

Lindberg:2023:MLD

[LNRW23] Julia Lindberg, Nathan Nicholson, Jose I. Rodriguez, and Zinan Wang. The maximum likelihood degree of sparse polynomial systems. SIAM Journal on Applied Algebra and Geometry, 7(1):159– 171, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1422550.

Le:2023:TCM

[LR23] Dinh V. Le and Tim Römer.

Theorems of Carathéodory, Minkowski–Weyl, and Gordan up to symmetry. SIAM Journal on Applied Algebra and Geometry, 7(1):291–310, ???? 2023. CODEN SJAABQ. ISSN 2470-

6566. URL https://epubs.
siam.org/doi/10.1137/22M148865X.

Lesnick:2020:QGI

[LRR20] Michael Lesnick, Raúl Rabadán, and Daniel I. S. Rosenbloom. Quantifying genetic innovation: Mathematical foundations for the topological study of reticulate evolution. SIAM Journal on Applied Algebra and Geometry, 4(1):141–184, ???? 2020. CODEN SJAABQ. ISSN 2470-6566.

Ladisch:2018:ELO

[LS18a] Frieder Ladisch and Achill Schürmann. Equivalence of lattice orbit polytopes. SIAM Journal on Applied Algebra and Geometry, 2(2):259–280, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Little:2018:CSS

[LS18b] John Little and Hal Schenck. Codes from surfaces with small Picard number. SIAM Journal on Applied Algebra and Geometry, 2(2):242–258, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Lairez:2019:NTM

[LS19] Pierre Lairez and Emre Can Sertöz. A numerical transcendental method in algebraic geometry: Computation of Picard groups and related invariants. SIAM Journal on Applied Algebra and Geometry, 3(4):559–584, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Lieblich:2020:THS

[LV20] Max Lieblich and Lucas Van Meter. Two Hilbert schemes in computer vision. SIAM Journal on Applied Algebra and Geometry, 4(2):297–321, ???? 2020. CODEN SJAABQ. ISSN 2470-6566.

Lesnick:2022:CMP

[LW22] Michael Lesnick and Matthew Wright. Computing minimal presentations and bigraded Betti numbers of 2parameter persistent homology. SIAM Journal on Applied Algebra and Geometry, 6(2):267– 298, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1388425.

Miolane:2018:TCT

[MHP18] Nina Miolane, Susan Holmes, and Xavier Pennec. Topologically constrained template estimation via Morse–Smale complexes controls its statistical consistency. SIAM Journal on Applied Algebra and Geometry, 2(2):348–375, ???? 2018. CO-DEN SJAABQ. ISSN 2470-6566.

Muller:2019:BFE

[MHR19] Stefan Müller, Josef Hofbauer, and Georg Regensburger. On the bijectivity of families of exponential/generalized polynomial maps. SIAM Journal on Applied Algebra and Geometry, 3(3):412–438, ???? 2019. CO-DEN SJAABQ. ISSN 2470-6566.

Micheli:2019:SPD

[Mic19] Giacomo Micheli. On the selection of polynomials for the DLP quasi-polynomial time algorithm for finite fields of small characteristic. SIAM Journal on Applied Algebra and Geometry, 3(2):256–265, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Monod:2019:TSS

[MKPGC19] Anthea Monod, Sara Kalisnik, Juan Ángel Patiño-Galindo, and Lorin Crawford. Tropical sufficient statistics for persistent homology. SIAM Journal on Applied Algebra and Geometry, 3 (2):337–371, ???? 2019. CO-DEN SJAABQ. ISSN 2470-6566.

Munthe-Kaas:2020:ICL

[MKSV20] Hans Z. Munthe-Kaas, Ari Stern, and Olivier Verdier. Invariant connections, Lie algebra actions, and foundations of numerical integration on manifolds. SIAM Journal on Applied Algebra and Geometry, 4 (1):49–68, ???? 2020. CODEN SJAABQ. ISSN 2470-6566.

Montufar:2017:DMK

[MM17] Guido Montúfar and Jason Morton. Dimension of marginals of Kronecker product models. SIAM Journal on Applied Algebra and Geometry, 1(1):126–151, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Michalek:2021:MLD

[MMW21] Mateusz Michałek, Leonid Monin, and Jarosław A. Wiśniewski. Maximum likelihood degree, complete quadrics, and C^{∗}- action. SIAM Journal on Applied Algebra and Geometry, 5 (1):60–85, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1335960.

McCleary:2022:EDP

[MP22] Alexander McCleary and Amit Patel. Edit distance and persistence diagrams over lattices. SIAM Journal on Applied Algebra and Geometry, 6(2):134– 155, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1373700.

Maxim:2020:EDD

[MRW20] Laurentiu G. Maxim, Jose I. Rodriguez, and Botong Wang. Euclidean distance degree of the multiview variety. SIAM Journal on Applied Algebra and Geometry, 4(1):28–48, ???? 2020. CODEN SJAABQ. ISSN 2470-6566.

Montufar:2022:SBN

[MRZ22] Guido Montúfar, Yue Ren, and Leon Zhang. Sharp bounds for the number of regions of maxout networks and vertices of Minkowski sums. SIAM Journal on Applied Algebra and Geometry, 6(4):618– 649, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1413699.

Meunier:2019:MVS

[MS19] Frédéric Meunier and Francis Edward Su. Multilabeled versions of Sperner’s and Fan’s lemmas and applications. SIAM Journal on Applied Algebra and Geometry, 3(3):391–411, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Malic:2023:CCP

[MS23] Goran Mali’c and Ileana Streinu. Computing circuit polynomials in the algebraic rigidity matroid. SIAM Journal on Applied Algebra and Geometry, 7(2):345– 385, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1437986.

Meyer:2022:ZST

[MSS22] Kyle Meyer, Ivan Soprunov, and Jenya Soprunova. F_{q}- zeros of sparse trivariate polynomials and toric 3-fold codes. SIAM Journal on Applied Algebra and Geometry, 6(3):432– 467, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1436890.

Neff:2019:EGM

[NFB19] Patrizio Neff, Andreas Fischle, and Lev Borisov. Explicit global minimization of the symmetrized Euclidean distance by a characterization of real matrices with symmetric square. SIAM Journal on Applied Algebra and Geometry, 3(1):31–43, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Nicklasson:2023:TNT

[Nic23] Lisa Nicklasson. Toric and non-toric Bayesian networks.

SIAM Journal on Applied Algebra and Geometry, 7(3):549– 566, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/22M1515690.

Nie:2017:STN

[Nie17] Jiawang Nie. Symmetric tensor nuclear norms. SIAM Journal on Applied Algebra and Geometry, 1(1):599–625, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Needham:2023:GAP

[NW23] Tom Needham and Thomas Weighill. Geometric averages of partitioned datasets. SIAM Journal on Applied Algebra and Geometry, 7(1):104– 132, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1436580.

Obayashi:2018:VOC

[Oba18] Ippei Obayashi. Volume-optimal cycle: Tightest representative cycle of a generator in persistent homology. SIAM Journal on Applied Algebra and Geometry, 2(4):508–534, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Ovchinnikov:2022:MEP

[OPPS22] Alexey Ovchinnikov, Anand Pillay, Gleb Pogudin, and Thomas Scanlon. Multiexperiment parameter identifiability of ODEs and model theory. SIAM Journal on Applied Algebra and Geometry, 6(3): 339–367, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1389845.

Oertel:2020:DFR

[OPW20] Timm Oertel, Joseph Paat, and Robert Weismantel. The distributions of functions related to parametric integer optimization. SIAM Journal on Applied Algebra and Geometry, 4(3):422–440, ???? 2020. CODEN SJAABQ. ISSN 2470-6566.

Ozyesil:2018:SCM

[ÖSS18] Onur Özyesil, Nir Sharon, and Amit Singer. Synchronization over Cartan motion groups via contraction. SIAM Journal on Applied Algebra and Geometry, 2(2):207–241, ???? 2018. CO-DEN SJAABQ. ISSN 2470-6566.

Polak:2022:SRO

[Pol22] Sven C. Polak. Symmetry reduction to optimize a graph-based polynomial from queueing theory. SIAM Journal on Applied Algebra and Geometry, 6(2): 243–266, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1413298.

Pryhuber:2022:ETV

[PST22] Andrew Pryhuber, Rainer Sinn, and Rekha R. Thomas. Existence of two view chiral reconstructions. SIAM Journal on Applied Algebra and Geometry, 6(1):41–76, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1381848.

Plaumann:2021:KTJ

[PSW21] Daniel Plaumann, Rainer Sinn, and Stephan Weis. Kippenhahn’s theorem for joint numerical ranges and quantum states. SIAM Journal on Applied Algebra and Geometry, 5 (1):86–113, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/19M1286578.

Ravagnani:2022:WNC

[Rav22]  Alberto Ravagnani.  Whit-

ney numbers of combinatorial geometries and higher-weight Dowling lattices. SIAM Journal on Applied Algebra and Geometry, 6(2):156–189, ???? 2022. CODEN SJAABQ. ISSN 2470-

6566. URL https://epubs.
siam.org/doi/10.1137/20M1382635.

Robeva:2021:MTS

[RS21] Elina Robeva and Jean-Baptiste Seby. Multi-trek separation in linear structural equation models. SIAM Journal on Applied Algebra and Geometry, 5(2): 278–303, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1316470.

Rodriguez:2017:MLD

[RW17] Jose Israel Rodriguez and Botong Wang. The maximum likelihood degree of mixtures of independence models. SIAM Journal on Applied Algebra and Geometry, 1(1):484–506, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Saunderson:2019:CPN

[Sau19] James Saunderson. Certifying polynomial nonnegativity via hyperbolic optimization. SIAM Journal on Applied Algebra and Geometry, 3(4):661–690, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Scheiderer:2018:SRC

[Sch18a] Claus Scheiderer. Semidefinite representation for convex hulls of real algebraic curves. SIAM Journal on Applied Algebra and Geometry, 2(1):1–25, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Scheiderer:2018:SS

[Sch18b] Claus Scheiderer. Spectrahedral shadows. SIAM Journal on Applied Algebra and Geometry, 2 (1):26–44, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Scheiderer:2021:SOC

[Sch21]  Claus Scheiderer.  Second-

order cone representation for convex sets in the plane. SIAM Journal on Applied Algebra and Geometry, 5(1):114– 139, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/20M133717X.

Shen:2017:SBR

[SG17] Li-Yong Shen and Ron Goldman. Strong μ-bases for rational tensor product surfaces and extraneous factors associated to bad base points and anomalies at infinity. SIAM Journal on Applied Algebra and Geometry, 1(1):328–351, ???? 2017. CO-DEN SJAABQ. ISSN 2470-6566.

Shankar:2019:CGD

[Sha19] Shiva Shankar. Corrigendum:

The Generic Degree of Autonomy. SIAM Journal on Applied Algebra and Geometry, 3 (1):172–174, ???? 2019. CO-DEN SJAABQ. ISSN 2470-6566. See [SR18].

Shitov:2018:CCC

[Shi18] Yaroslav Shitov. A counterexample to Comon’s conjecture. SIAM Journal on Applied Algebra and Geometry, 2(3):428–443, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Shitov:2021:WRX

[Shi21] Yaroslav Shitov. The Waring rank of the 3 x 3 permanent. SIAM Journal on Applied Algebra and Geometry, 5(4):701– 714, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1349254.

Shitov:2022:SCS

[Shi22] Yaroslav N. Shitov. The subrank of a complex symmetric tensor can exceed its symmetric subrank. SIAM Journal on Applied Algebra and Geometry, 6(4): 531–535, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1465494.

Shankar:2018:GDA

[SR18] Shiva Shankar and Paula Rocha. The generic degree of autonomy. SIAM Journal on Applied Algebra and Geometry, 2(3):410–427, ???? 2018. CODEN SJAABQ. ISSN 2470-6566. See corrigendum [Sha19].

Stange:2021:AAS

[Sta21] Katherine E. Stange. Algebraic aspects of solving Ring-LWE, including ring-based improvements in the Blum– Kalai–Wasserman algorithm. SIAM Journal on Applied Algebra and Geometry, 5(2):366– 387, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/19M1280442.

Sturmfels:2017:MEC

[Stu17] Bernd Sturmfels. Message from the Editor-in-Chief. SIAM Journal on Applied Algebra and Geometry, 1(1):1, ???? 2017. CO-DEN SJAABQ. ISSN 2470-6566.

Tokcan:2021:AMT

[TGND21] Neriman Tokcan, Jonathan Gryak, Kayvan Najarian, and Harm Derksen. Algebraic methods for tensor data. SIAM Journal on Applied Algebra and Geometry, 5(1):1–27, ???? 2021. CODEN SJAABQ. ISSN 2470-

6566. URL https://epubs.
siam.org/doi/10.1137/19M1272494.

Tran:2021:FMB

[Tra21] Ngoc Mai Tran. The finite matroid-based valuation conjecture is false. SIAM Journal on Applied Algebra and Geometry, 5(3):506–525, ???? 2021. CODEN SJAABQ. ISSN 2470-

6566. URL https://epubs.
siam.org/doi/10.1137/19M1304295.

Turatti:2022:TDT

[Tur22] Ettore Turatti. On tensors that are determined by their singular tuples. SIAM Journal on Applied Algebra and Geometry, 6 (2):319–338, ???? 2022. CO-DEN SJAABQ. ISSN 2470-

6566. URL https://epubs.
siam.org/doi/10.1137/21M1412980.

Volcic:2019:SNP

[Vol19] Jurij Volcic. Stable noncommutative polynomials and their determinantal representations. SIAM Journal on Applied Algebra and Geometry, 3(1):152–171, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Wang:2022:NPC

[Wan22] Jie Wang. Nonnegative polynomials and circuit polynomials. SIAM Journal on Applied Algebra and Geometry, 6(2): 111–133, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1313969.

Ye:2017:CCE

[YL17] Ke Ye and Lek-Heng Lim. Cohomology of cryo-electron microscopy. SIAM Journal on Applied Algebra and Geometry, 1 (1):507–535, ???? 2017. CO-DEN SJAABQ. ISSN 2470-6566.

Yu:2021:OCA

[YMV21] Haizi Yu, Igor Mineyev, and Lav R. Varshney. Orbit computation for atomically generated subgroups of isometries of Z^{n}. SIAM Journal on Applied Algebra and Geometry, 5(3): 479–505, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/20M1375127.